How much principal have you paid by year 10 on a 30year mortgage for $175,000, with a 5% interest rate, paid monthly.

On paper with work please
Answer #1

Monthly payment=175000*(5%/12)/(1-1/(1+5%/12)^(12*30))=939.4378403

Loan oustanding after 10 years=175000*(1+5%/12)^(12*10)-939.4378403/(5%/12)*((1+5%/12)^(12*10)-1)=142348.6129

Principal paid=175000-142348.6129=32651.3871